# Mamba-FSCIL: Dynamic Adaptation with Selective State Space Model for Few-Shot Class-Incremental Learning

### v1

@@ 5.4 Analysis of Feature Embeddings
We analyze feature embeddings using t-SNE visualizations for both base and novel classes from the test set of the CIFAR-100 dataset, comparing the strong baseline NC-FSCIL with our Mamba-FSCIL. Fig. 3-(a) and Fig. 3-(b) show the results of some base classes during the first and last sessions. It is shown that the base-class features of the baseline method are much scattered in the last session. As a comparison, Mamba-FSCIL demonstrates consistent intra-class compactness and distinct inter-class separation, indicating enhanced stability for base classes that better maintain their performance in incremental training. Fig. 3-(c) and Fig. 3-(d) depict the results of some novel classes in the 6-th and 8-th sessions. In both sessions, Mamba-FSCIL shows significantly more compact clusters with clearer margins among classes, highlighting its capacity to accommodating new classes. These results explain Mamba-FSCIL’s superior performance in the benchmarks.
5.5 Components in Class-sensitive Selective Scan
We investigate the impacts of the suppression and separation losses introduced in Sec. 4.3. First, we measure the average cosine similarity between the final representations and the classifier prototypes across all base classes in Fig. 4, defined as:
$\frac{1}{\sum_{k=1}^{K}N_{k}}\sum_{k=1}^{K}\sum_{i=1}^{N_{k}}\cos\angle({\bm{\mu}}_{k,i},\mathbf{w}_{k}),$ (23)
where ${\bm{\mu}}_{k,i}$ represents the feature vector of sample $i$ in class $k$, $N_{k}$ is the number of samples in class $k$, $\mathbf{w}_{k}$ is the classifier prototype for class $k$, and $K$ is the number of base classes. To assess the separability between base and novel classes, we calculate the average cosine similarity for any possible pairing of base and novel classes feature vectors as:
$\frac{1}{N_{\text{base}}N_{\text{novel}}}\sum_{i=1}^{N_{\text{base}}}\sum_{j=1}^{N_{\text{novel}}}\cos\angle({\bm{\mu}}_{\text{base},i},{\bm{\mu}}_{\text{novel},j}),$ (24)
where ${\bm{\mu}}_{\text{base},i}$ and ${\bm{\mu}}_{\text{novel},j}$ are feature vectors of the base class sample $i$ and novel class sample $j$, respectively.
Fig. 4 shows that the incorporation of suppression loss ${\mathcal{L}}_{\text{supp}}$ helps to improve the base-class alignment between features and classifier prototypes in late sessions, facilitating the maintenance of base-class performance. When the separation loss ${\mathcal{L}}_{\text{sep}}$ is further added, the similarities of features between the base and novel classes are significantly reduced, which leaves sufficient feature space for incremental training sessions, promoting the novel-class accuracy.
@@ 5.5 Components in Class-sensitive Selective Scan
Figure 4: Impact of suppression and separation losses on the alignment between base-class feature and classifier and the separability between base and novel classes in MiniImageNet’s test set: (a) average cosine similarity of base class features with their corresponding classifiers; (b) average cosine similarity between base-class and novel-class features.
5.6 Comparison with Dynamic Adaptation Method
We compare our Mamba-FSCIL with DSN [68] in Fig. 5, particularly in terms of parameter efficiency and session-wise performance. Both methods use the same ResNet18 backbone. As shown in Fig. 5-(a), Unlike DSN, which continually increases its parameter amount to learn new classes, Mamba-FSCIL armed with a dual selective SSM projector dynamically adjusts its parameters without expanding the parameter space. In addition, Fig. 5-(b) illustrates that Mamba-FSCIL consistently outperforms DSN in all sessions, demonstrating Mamba-FSCIL’s superior ability to retain old knowledge while learning new classes.
6 Conclusion
In this study, we propose the Mamba-FSCIL framework, leveraging selective state space models to address the challenges of FSCIL. Mamba-FSCIL utilizes a dual selective SSM projector to achieve dynamic adaptation without the need to expand the model’s parameter space continually. The dual design dynamically adjusts its parameters to integrate new classes while preserving the integrity of previously learned information. Furthermore, the incorporation of the class-sensitive selective scan mechanism, composed of the suppression loss and the separation loss, guides the dynamic adaptation process deliberately, enhancing both stability and adaptability in incremental sessions’ training. Empirical results demonstrate that Mamba-FSCIL outperforms the state-of-the-art static and dynamic methods across three benchmark datasets.

### v2

@@ 5.7 Analysis of Feature Embeddings
We analyzed feature embeddings using t-SNE visualizations for both base and novel classes from the test set of the CIFAR-100 dataset, comparing the strong baseline NC-FSCIL with our Mamba-FSCIL. Fig. 8 (a) and Fig. 8 (b) show the results of some base classes during the first and last sessions. It is shown that the base-class features of the baseline method are much scattered in the last session. As a comparison, Mamba-FSCIL demonstrates consistent intra-class compactness and distinct inter-class separation, indicating enhanced stability for base classes that better maintain their performance in incremental training. Fig. 8 (c) and Fig. 8 (d) depict the results of some novel classes in the 6-th and 8-th sessions. In both sessions, Mamba-FSCIL shows significantly more compact clusters with clearer margins among classes, highlighting its capacity to accommodate new classes. These results explain Mamba-FSCIL’s superior performance in the benchmarks.
6 Conclusion and Future Works
In this study, we propose the Mamba-FSCIL framework, leveraging selective state space models to address the challenges of FSCIL. Mamba-FSCIL utilizes a dual selective SSM projector to achieve dynamic adaptation without the need to continually expand the parameter space of the model. The dual design dynamically adjusts its parameters to integrate new classes while preserving the integrity of previously learned information. Furthermore, the incorporation of the class-sensitive selective scan mechanism, composed of the suppression loss and the separation loss, guides the dynamic adaptation process deliberately, enhancing both stability and adaptability in incremental sessions’ training. Empirical results demonstrate that Mamba-FSCIL outperforms the state-of-the-art static and dynamic methods across three benchmark datasets.
Future work for Mamba-FSCIL involves extending it to address highly correlated tasks like online class-incremental and cross-domain incremental learning. We also aim to adapt Mamba-FSCIL for multimodal scenarios, enhancing its effectiveness in dynamic and interactive environments. Additionally, exploring its application in embodied intelligence is also a key direction for future research.